Nova Inc. has liabilities of 10 due in 1, 4, and 7 years. What asset incomes must the company arrange for in years 1 and 6 to immunize their cash flow, assuming an annual interest rate of 10% on all transactions?
Answer #1

x/1.1+x/1.1^6=10/1.1+10/1.1^4+10/1.1^7
=>x=(10/1.1+10/1.1^4+10/1.1^7)/(1/1.1+1/1.1^6)
=>x=14.2868670